A box measures 90 mm, 31.5 mm by 6.3 mm. State its volume with the proper number of significant figures with correct units.
1
Expert's answer
2020-10-09T13:46:42-0400

A box measures 90 mm, 31.5 mm by 6.3 mm.


So , Volume is = ( 90 * 31.5 * 6.3 ) mm^3 . = 17860.5 mm^3 . = 1.8 * 10 ^4 mm ^3 . answer